A car is traveling at a speed of 53 feet per second. (a) What is its speed in kilometers per hour? km/h (b) Is it exceeding the 40 mile per hour speed limit? O Yes No
Answer #1

(a) 1 km = 3280.84 feet

1 hr = 3600 s

1 feet/s = 3600/3280.84 = 1.097 km/hr

53 feet /s = 53*1.097 km/hr = 58.14 km/hr

(b) 1 km = 0.621 mile

58.14 km /hr = 58.14*0.621 miles/hr = 36.10 miles/hr

No